OpenStudy (anonymous):

A circle has an area of 441π cm2. What is the circumference? a.21π cm b.42π cm c.42 cm d.21 cm

OpenStudy (anonymous):

A=pi*r^2 C=2*pi*r

OpenStudy (anonymous):

Yeah I know but I dont know how to get the circumference from the area

OpenStudy (anonymous):

...

OpenStudy (anonymous):

since you have the area of a circle, then you can find the radius of the circle once you determine the radius of the circle, you can then use it to calculate the circumference of the circle

OpenStudy (anonymous):

okay so 441= 3.14r^2